# A Radio Map Approach for Reduced Pilot CSI Tracking in Massive MIMO Networks

### v3

@@ IV-C An Iterative Framework for Radio Map Construction
      Estimate the channel covariance $\hat{\mathbf{C}}^{(l)}(\mathbf{x}_{i})$ using $\mathcal{Y}_{T}$, $\hat{\mathcal{P}}_{T}^{(l)}$, and (37), and construct the radio map ^$\hat{\mathcal{M}}^{(l)}=\{(\mathbf{x},\hat{\mathbf{C}}^{(l)}({\bf x})):\mathbf{x}\in\mathcal{X}\}$.
@@
    Output ^$\hat{\mathcal{M}}^{(l)}$ and $\hat{\mathcal{P}}_{T}^{(l)}$.
@@ V-A Environment Setup and Scenarios
The experiments are conducted on a MIMO CSI database generated by Wireless Insite [^1] http://www.remcom.com/wireless-insite. which provides 3D ray-tracing for the analysis of site-specific radio wave propagation. The tested city topology is a $710$-meter $\times$ $740$-meter area from San Francisco, America. Seven BSs are deployed on the top of some buildings as shown in Fig. 4(a). The MIMO antenna of each BS consists of $N_{\mathrm{t}}=64$ antenna arrays as shown in Fig. 4(b). By simulating random walks along roads in the map, the radio database is established, and it contains $19331$ user locations with $19331\times 7$ MIMO channels.
@@
Note that Doppler shift considerations are not necessary in our experimental setup. This is because the primary purpose of CSI estimation in the proposed scheme is to enable MIMO beamforming at the BS. Specifically, it aims to find a beamforming vector ${\bf b}$ that maximizes the signal strength $|{\bf b}^{\text{H}}{\bf h}e^{-j\theta}|$, where the phase offset $\theta$ due to the Doppler effect does not affect the signal strength. During beamformed data transmission, additional pilot symbols are still required to compensate the phase offset $\theta$. However, this aspect is beyond the scope of this work, which centers on reducing pilot overhead for CSI tracking to enable effective beamforming in mobile scenarios.
@@ V-B Baseline Schemes
The following metrics are used to evaluate the proposed algorithms: 1) Localization error: Given the true user position $\mathbf{p}_{t}$ and the estimated user position $\hat{\mathbf{p}}$, the localization error is defined as $\frac{1}{T}\sum_{t=1}^{T}\|{\bf p}_{t}-\hat{{\bf p}}_{t}\|_{2}.$ 2) Channel capacity: Using the maximum ratio combining (MRC) scheme, the channel capacity is expressed as $f(\mathbf{b},{\bf h})=\log_{2}(1+|\mathbf{b}^{\text{H}}{\bf h}|^{2}/\sigma_{\text{n}}^{2}),$ where ${\bf h}$ is the true CSI and $\mathbf{b}=\hat{\mathbf{h}}/\|\hat{\mathbf{h}}\|_{2}$ is the beamforming vector, in which $\hat{\mathbf{h}}$ is the estimated CSI. 3) Channel capacity efficiency ratio: This is defined as the ratio of the maximum capacity achieved with $\hat{\mathbf{h}}$ relative to that with perfect CSI ${\bf h}$. 4) Normalized $L_{2}$-norm estimation error of channel covariance: This is defined as $\frac{1}{|\mathcal{X}|}\sum_{\mathbf{x}\in\mathcal{X}}\|\hat{\mathbf{C}}(\mathbf{x})-\mathbf{C}(\mathbf{x})\|_{2}/\|\mathbf{C}(\mathbf{x})\|_{2}$.
@@ V-C Radio-Map-Embedded CSI Tracking
We first evaluate the proposed radio-map-embedded CSI tracking scheme. A perfect radio map $\mathcal{M}$ for each BS is constructed by discretizing the area of interest and calculating $\mathbf{C}(\mathbf{x}_{i})$ using true channels in each grid cell centered at $\mathbf{x}_{i}$. Fig. 5 illustrates the CSI and user position tracking performance of Algorithm 1. Specifically, Fig. 5(a) shows the channel capacity along a trajectory transitioning through NLOS, LOS, and NLOS regions to BS 7. The proposed scheme outperforms other baseline schemes, ensuring smooth tracking during the NLOSLOS transition. Fig. 5(b) reveals that the channel capacity efficiency ratio of the proposed scheme remains above $90\%$ for an LOS trajectory to BS 1. The tracking performance of the conventional KF scheme is unstable due to the lack of prior CSI statistics, while the CAM-aided scheme achieves above $78\%$ of capacity over that of perfect CSI. Fig. 5(c) highlights that localization errors along the same trajectory in Fig. 5(a) stay below $26$ meters, where BS 1 operates under LOS conditions and BS 6 under NLOS. Additionally, the presence of multiple BSs contributes to more stable performance and reduced localization errors, with values consistently below $14$ meters. For multi-BS cooperation, each BS independently computes the position distribution using (21), and then, these distributions are combined multiplicatively across BSs to form a joint posterior. Subsequently, the position ${\bf p}_{t}\in\mathcal{X}$ that maximizes this posterior is selected according to (22).
Fig. 6 shows the channel capacity efficiency ratio versus SNR when $7$ BSs serve the user. This is merely an illustrative example to demonstrate the potential performance gains achievable with multiple BSs. The proposed scheme achieves over $97\%$ of the capacity attained with perfect CSI at an SNR of $20$ dB. Additionally, the proposed scheme nearly matches the performance of the genius-aided scheme, which assumes perfect user positions. The one-shot CAM scheme, which utilizes only the current measurement and $32$ pilots ($M=32$), achieves approximately $87\%$ of the ideal capacity at an SNR of $30$ dB. The KF scheme performs well at high SNR but achieves only about $36\%$ of the ideal capacity at an SNR of $10$ dB. The AR-based and LSTM-based methods perform poorly due to their dependence on highly accurate historical channel data and restrictive assumptions about stationary spatial distributions. The LS estimator consistently underperforms other baselines, because it fails to exploit spatial and temporal priors and relies solely on a one-shot observation for CSI estimation.
@@
Table I shows the impact of adaptive sensing matrices in the proposed CSI tracking scheme. It is observed that adaptive sensing provides additional performance gains, especially in low-SNR scenarios. For example, at $10$ dB SNR under NLOS conditions to BS 6, the proposed method with adaptive ${\bf A}_{t}$ achieves a channel capacity efficiency ratio of $55\%$, representing a $14\%$ improvement over the same scheme with random ${\bf A}_{t}$. In addition, both variants of the proposed scheme with adaptive and random ${\bf A}_{t}$ achieves over $91\%$ of the idea capacity under LOS conditions to BS 1, demonstrating the robustness of the proposed radio-map-embedded CSI tracking scheme, thanks to the strong spatial prior provided by the radio map.
@@ V-D Radio Map Construction
In this subsection, we construct radio maps for each BS from sparse and unlabeled pilot sequences in an offline manner. This is a one-time process performed prior to deployment. Fig. 8(a) shows an example of the estimated user trajectory obtained by solving Problem (34). The estimated trajectory matches the ground truth well, even though the coarse user locations are scattered over an area hundreds of meters wide. Numerical results indicate a localization error of less than $6$ meters when BS $1$ serves the user. Fig. 8(b) confirms the convergence of Algorithm 2 with localization errors staying stable as the number of iterations increases.
@@
Fig. 9 and Fig. 10 illustrate the CSI tracking performance for BS $1$ under LOS conditions and BS $6$ under NLOS, based on the radio maps that are reconstructed using $280,000$ and $420,000$ CSI measurements, respectively with $M=16$. Using reconstructed radio maps, the proposed CSI tracking scheme with one pilot at each time slot and a total number of $100$ pilots achieves above $88\%$ of the ideal capacity for BS 1, and the proposed scheme outperforms other baseline schemes in both low-SNR and high-SNR scenarios for BS $6$.
Additionally, the memory usage of the proposed radio map is assessed. The radio map covers an area of $740$ m×$\times$$710$ m with a grid resolution of $5$ m, resulting in $21,016$ grids. Each grid stores its center coordinates as double-precision values and a $64\times 64$ complex covariance matrix in single precision. The theoretical storage requirement is approximately $657$ MB, and the actual uncompressed size is about $659$ MB. This demonstrates the feasibility of the proposed method, as it achieves a reasonable memory size while maintaining the necessary precision for the radio map.
@@ VI Conclusion
This paper addresses the challenges of CSI tracking in massive MIMO systems, particularly in the absence of stationary CSI statistics and precise location information. By integrating radio maps with an adaptive SKF framework, we track the CSI and the location using an extremely sparse pilot sequence without location labels. For radio map construction, the joint estimation of location sequences and channel covariance matrices is facilitated by an HMM, and an unbiased channel covariance estimator is found with the estimation error analyzed. Numerical results show that the proposed radio-map-embedded CSI tracking scheme delivers over $97\%$ of the capacity of perfect CSI with reduced pilots at a $20$ dB SNR under LOS, while a conventional KF can only achieve $76\%$. In addition, the proposed algorithm reduces localization errors from $30$ meters from the prior to $6$ meters for radio map construction in our framework.
@@
$=\prod_{t=1}^{T}p(\mathbf{y}_{\text{$t$}}|\mathbf{h}_{t})\times\prod_{t=2}^{T}p(\mathbf{h}_{\text{$t$}}|\mathbf{h}_{t-1},{\bf p}_{t})p(\mathbf{h}_{1}),$ (42)
@@
$p(\mathcal{H}_{T},\mathcal{P}_{T})$ =$\prod_{t=2}^{T}p({\bf h}_{t}|{\bf h}_{t-1},{\bf p}_{t})\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})p({\bf h}_{1})\mathbb{P}(\mathbf{p}_{1}).$ (44)
@@
×$\quad\times\prod_{t=2}^{T}\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})\mathbb{P}(\mathbf{p}_{1}),$ (46)
@@
×$\times{\bf A}_{t}\mathbf{Q}_{t|t-1})(\mathbf{Q}_{t|t-1}^{-1}+\mathbf{A}_{t}^{\text{H}}(\sigma_{\mathrm{n}}^{2}\mathbf{I})^{-1}\mathbf{A}_{t})$
@@
$\mathbf{Q}_{t|t-1}=\mathbf{W}_{t-1}\boldsymbol{\Lambda}_{t-1}\mathbf{W}_{t-1}^{\mathrm{H}},$ (48)
@@
Thus, $|\mathbf{B}_{t}(\mathbf{I}+\sigma_{\text{n}}^{-2}\boldsymbol{\Lambda}_{t-1})\mathbf{B}_{t}^{\mathrm{H}}|$ reaches its maximum as $\prod_{i=1}^{M}(1+\sigma_{\text{n}}^{2}\lambda_{t-1,i}),$ where $\lambda_{t-1,i}$ for $i=1,2,\dots,M$ are the top $M$ eigenvalues of $\mathbf{Q}_{t|t-1}$. In this case, $\mathbf{A}_{t}$ is given by $\mathbf{A}_{t}=[\mathbf{w}_{t-1,1},\mathbf{w}_{t-1,2},\dots,\mathbf{w}_{t-1,M}]^{\text{H}}$ such that $\mathbf{B}_{t}=\mathbf{A}_{t}\mathbf{W}_{t-1}$ is a diagonal matrix with its first $M$ diagonal elements equal to $1$ and other elements equal to 00.
@@ Appendix D Proof of Proposition 4
$=\frac{N_{\mathrm{t}}^{2}}{M^{2}}(\mathbb{E}\{\boldsymbol{\Phi}\mathbf{h}(\boldsymbol{\Phi}\mathbf{h})^{\text{H}}|\mathbf{h}\}+\mathbb{E}\{\mathbf{A}^{\text{H}}\mathbf{n}\mathbf{n}^{\mathrm{H}}\mathbf{A}\}),$
@@
Second, the following fact is adopted from [31, Fact 1].
@@
$\boldsymbol{\Phi}\mathbf{x}\overset{\mathrm{d}}{=}\omega\mathbf{x}+\sqrt{\omega-\omega^{2}}\|\mathbf{x}\|\mathbf{W}\alpha,$
@@
$-\frac{\sigma_{\text{n}}^{2}(M\bar{N}_{\mathrm{t}}\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i})-N_{\mathrm{t}}(N_{\mathrm{t}}-M)\text{tr}(\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i}))\mathbf{I})}{M(N_{\mathrm{t}}M+N_{\mathrm{t}}-2)},$ (54)

### v5

@@ IV-C An Iterative Framework for Radio Map Construction
      Estimate the channel covariance $\hat{\mathbf{C}}^{(l)}(\mathbf{x}_{i})$ using $\mathcal{Y}_{T}$, $\hat{\mathcal{P}}_{T}^{(l)}$, and (37), and construct the radio map $\hat{\mathcal{M}}^{(l)}=\{(\mathbf{x},\hat{\mathbf{C}}^{(l)}({\bf x})):\mathbf{x}\in\mathcal{X}\}$.
@@
    Output $\hat{\mathcal{M}}^{(l)}$ and $\hat{\mathcal{P}}_{T}^{(l)}$.
@@ V-A Environment Setup and Scenarios
The experiments are conducted on a MIMO CSI database generated by Wireless Insite [33] which provides 3D ray-tracing for the analysis of site-specific radio wave propagation. The tested city topology is a $710$-meter $\times$ $740$-meter area from San Francisco, America. Seven BSs are deployed on the top of some buildings as shown in Fig. 4(a). The MIMO antenna of each BS consists of $N_{\mathrm{t}}=64$ antenna arrays as shown in Fig. 4(b). By simulating random walks along roads in the map, the radio database is established, and it contains $19331$ user locations with $19331\times 7$ MIMO channels.
@@
Note that Doppler shift considerations are not necessary in our experimental setup. This is because the primary purpose of CSI estimation in the proposed scheme is to enable MIMO beamforming at the BS. Specifically, it aims to find a beamforming vector $\bf b$} that maximizes the signal strength $|{\bf b}^{\text{H}}{\bf h}e^{-j\theta}|$, where the phase offset $\theta$ due to the Doppler effect does not affect the signal strength. During beamformed data transmission, additional pilot observations are still required to compensate for the phase offset $\theta$. However, this aspect is beyond the scope of this work, which centers on reducing pilot overhead for CSI tracking to enable effective beamforming in mobile scenarios.
@@ V-B Baseline Schemes
The following metrics are used to evaluate the proposed algorithms: 1) Localization error: Given the true user position $\mathbf{p}_{t}$ and the estimated user position $\hat{\mathbf{p}}$, the localization error is defined as $\frac{1}{T}\sum_{t=1}^{T}\|{\bf p}_{t}-\hat{{\bf p}}_{t}\|_{2}.$ 2) Channel capacity: Using the maximum ratio combining (MRC) scheme, the channel capacity is expressed as $f(\mathbf{b},{\bf h})=\log_{2}(1+|\mathbf{b}^{\text{H}}{\bf h}|^{2}/\sigma_{\text{n}}^{2}),$ where $\bf h$} is the true CSI and $\mathbf{b}=\hat{\mathbf{h}}/\|\hat{\mathbf{h}}\|_{2}$ is the beamforming vector, in which $\hat{\mathbf{h}}$ is the estimated CSI. 3) Channel capacity efficiency ratio: This is defined as the ratio of the maximum capacity achieved with $\hat{\mathbf{h}}$ relative to that with perfect CSI $\bf h$}. 4) Normalized $L_{2}$-norm estimation error of channel covariance: This is defined as $\frac{1}{|\mathcal{X}|}\sum_{\mathbf{x}\in\mathcal{X}}\|\hat{\mathbf{C}}(\mathbf{x})-\mathbf{C}(\mathbf{x})\|_{2}/\|\mathbf{C}(\mathbf{x})\|_{2}$. 5) Normalized projection error of channel covariance: This is defined as $\|\mathbf{C}-\hat{\mathbf{C}}(\hat{\mathbf{C}}^{\mathrm{H}}\hat{\mathbf{C}})^{-1}\hat{\mathbf{C}}^{\mathrm{H}}\mathbf{C}\|_{\mathrm{F}}/\|\mathbf{C}\|_{\mathrm{F}}$, which quantifies the mismatch between the true and estimated signal subspaces.
@@ V-C Radio-Map-Embedded CSI Tracking
We first evaluate the proposed radio-map-embedded CSI tracking scheme. A perfect radio map $\mathcal{M}$ for each BS is constructed by discretizing the area of interest and calculating $\mathbf{C}(\mathbf{x}_{i})$ using true channels in each grid cell centered at $\mathbf{x}_{i}$. Fig. 5 illustrates the CSI and user position tracking performance of Algorithm 1. Specifically, Fig. 5(a) shows the channel capacity along a trajectory transitioning through NLOS, LOS, and NLOS regions to BS 7. The proposed scheme outperforms other baseline schemes, ensuring smooth tracking during the NLOS-LOS transition. Fig. 5(b) reveals that the channel capacity efficiency ratio of the proposed scheme remains above $90\%$ for an LOS trajectory to BS 1. The tracking performance of the conventional KF scheme is unstable due to the lack of prior CSI statistics, while the CAM-aided scheme achieves above $78\%$ of capacity over that of perfect CSI. Fig. 5(c) highlights that localization errors along the same trajectory in Fig. 5(a) stay below $26$ meters, where BS 1 operates under LOS conditions and BS 6 under NLOS. Additionally, the presence of multiple BSs contributes to more stable performance and reduced localization errors, with values consistently below $14$ meters. For multi-BS cooperation, each BS independently computes the position distribution using (21), and then, these distributions are combined multiplicatively across BSs to form a joint posterior. Subsequently, the position ${\bf p}_{t}\in\mathcal{X}$ that maximizes this posterior is selected according to (22).
Fig. 6 shows the channel capacity efficiency ratio versus SNR when $7$ BSs serve the user. This is merely an illustrative example to demonstrate the potential performance gains achievable with multiple BSs. The proposed scheme achieves over $97\%$ of the capacity attained with perfect CSI at an SNR of $20$ dB. Additionally, the proposed scheme nearly matches the performance of the genius-aided scheme, which assumes perfect user positions. The one-shot CAM scheme, which utilizes only the current measurement and $32$ pilot observations ($M=32$), achieves approximately $87\%$ of the ideal capacity at an SNR of $30$ dB. The KF scheme performs well at high SNR but achieves only about $36\%$ of the ideal capacity at an SNR of $10$ dB. The AR-based and LSTM-based methods perform poorly due to their dependence on highly accurate historical channel data and restrictive assumptions about stationary spatial distributions. The LS estimator consistently underperforms other baselines, because it fails to exploit spatial and temporal priors and relies solely on a one-shot observation for CSI estimation.
@@
Table I shows the impact of adaptive sensing matrices in the proposed CSI tracking scheme. It is observed that adaptive sensing provides additional performance gains, especially in low-SNR scenarios. For example, at $10$ dB SNR under NLOS conditions to BS 6, the proposed method with adaptive ${\bf A}_{t}$ achieves a channel capacity efficiency ratio of $55\%$, representing a $14\%$ improvement over the same scheme with random ${\bf A}_{t}$. In addition, both variants of the proposed scheme with adaptive and random ${\bf A}_{t}$ achieve over $91\%$ of the ideal capacity under LOS conditions to BS 1, demonstrating the robustness of the proposed radio-map-embedded CSI tracking scheme, thanks to the strong spatial prior provided by the radio map.
@@ V-D Radio Map Construction
In this subsection, we construct radio maps for each BS from sparse and unlabeled pilot observation sequences in an offline manner. This is a one-time process performed prior to deployment. Fig. 8(a) shows an example of the estimated user trajectory obtained by solving Problem (34). The estimated trajectory matches the ground truth well, even though the coarse user locations are scattered over an area hundreds of meters wide. Numerical results indicate a localization error of less than $6$ meters when BS $1$ serves the user. Fig. 8(b) confirms the convergence of Algorithm 2 with localization errors staying stable as the number of iterations increases.
@@
Table V: Normalized projection error of channel covariance versus number of measurements $T_{\mathrm{total}}$ and dimension $M$
$T_{\mathrm{total}}$ $35000$ $70000$ $105000$ $210000$ $280000$
$M=1$, BS 1 (LOS) $0.227$ $0.227$ $0.203$ $0.139$ $0.131$
$M=4$, BS 1 (LOS) $0.235$ $0.207$ $0.199$ $0.131$ $0.131$
$M=16$, BS 1 (LOS) $0.211$ $0.131$ $0.092$ $0.064$ $0.052$
$M=16$, BS 6 (NLOS) $0.214$ $0.202$ $0.202$ $0.198$ $0.190$
@@
Table V further illustrates the normalized projection error of channel covariance versus the number of measurements $T_{\mathrm{total}}$ and dimension $M$. The results demonstrate that increased measurements or dimension of pilot observations consistently reduce the normalized projection error for both BS 1 and BS 6. However, BS 1, operating under LOS conditions, achieves a significantly lower normalized projection error with sufficient measurements. Specifically, BS 1 reaches a normalized projection error of $0.211$ with $35,000$ measurements and $0.052$ with $280,000$ measurements. In contrast, BS 6, operating under NLOS conditions, only achieves a normalized projection error of $0.190$ even with $280,000$ measurements. These findings indicate that radio map construction under NLOS conditions presents greater challenges than under LOS conditions due to rich multipath effects.
Fig. 9 and Fig. 10 illustrate the CSI tracking performance for BS $1$ under LOS conditions and BS $6$ under NLOS, based on the radio maps that are reconstructed using $280,000$ and $420,000$ CSI measurements, respectively with $M=16$. Using reconstructed radio maps, the proposed CSI tracking scheme with one pilot observation at each time slot and a total number of $100$ pilot observations achieves above $88\%$ of the ideal capacity for BS 1, and the proposed scheme outperforms other baseline schemes in both low-SNR and high-SNR scenarios for BS $6$.
[…]
@@ VI Conclusion
This paper addresses the challenges of CSI tracking in massive MIMO systems, particularly in the absence of stationary CSI statistics and precise location information. By integrating radio maps with an adaptive SKF framework, we track the CSI and the location using an extremely sparse pilot observation sequence without location labels. For radio map construction, the joint estimation of location sequences and channel covariance matrices is facilitated by an HMM, and an unbiased channel covariance estimator is found with the estimation error analyzed. Numerical results show that the proposed radio-map-embedded CSI tracking scheme delivers over $97\%$ of the capacity of perfect CSI with reduced pilot observations at a $20$ dB SNR under LOS, while a conventional KF can only achieve $76\%$. In addition, the proposed algorithm reduces localization errors from $30$ meters from the prior to $6$ meters for radio map construction in our framework.
@@
$=\prod_{t=1}^{T}p(\mathbf{y}_{\text{$t$}}|\mathbf{h}_{t})\times\prod_{t=2}^{T}p(\mathbf{h}_{\text{$t$}}|\mathbf{h}_{t-1},{\bf p}_{t})p(\mathbf{h}_{1})$ (42)
@@
$p(\mathcal{H}_{T},\mathcal{P}_{T})$ $\text{$=$}\prod_{t=2}^{T}p({\bf h}_{t}|{\bf h}_{t-1},{\bf p}_{t})\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})p({\bf h}_{1})\mathbb{P}(\mathbf{p}_{1}).$ (44)
@@
$\quad\times\prod_{t=2}^{T}\mathbb{P}(\mathbf{p}_{t}|\mathbf{p}_{t-1})\mathbb{P}(\mathbf{p}_{1})$ (46)
@@
$\times{\bf A}_{t}\mathbf{Q}_{t|t-1})(\mathbf{Q}_{t|t-1}^{-1}+\mathbf{A}_{t}^{\text{H}}(\sigma_{\mathrm{n}}^{2}\mathbf{I})^{-1}\mathbf{A}_{t})$
@@
$\mathbf{Q}_{t|t-1}=\mathbf{W}_{t-1}\boldsymbol{\Lambda}_{t-1}\mathbf{W}_{t-1}^{\mathrm{H}}$ (48)
@@
Thus, $|\mathbf{B}_{t}(\mathbf{I}+\sigma_{\text{n}}^{-2}\boldsymbol{\Lambda}_{t-1})\mathbf{B}_{t}^{\mathrm{H}}|$ reaches its maximum as $\prod_{i=1}^{M}(1+\sigma_{\text{n}}^{2}\lambda_{t-1,i}),$ where $\lambda_{t-1,i}$ for $i=1,2,\dots,M$ are the top $M$ eigenvalues of $\mathbf{Q}_{t|t-1}$. In this case, $\mathbf{A}_{t}$ is given by $\mathbf{A}_{t}=[\mathbf{w}_{t-1,1},\mathbf{w}_{t-1,2},\dots,\mathbf{w}_{t-1,M}]^{\text{H}}$ such that $\mathbf{B}_{t}=\mathbf{A}_{t}\mathbf{W}_{t-1}$ is a diagonal matrix with its first $M$ diagonal elements equal to $1$ and other elements equal to 0.
@@ Appendix D Proof of Proposition 4
$=\frac{N_{\mathrm{t}}^{2}}{M^{2}}(\mathbb{E}\{\boldsymbol{\Phi}\mathbf{h}(\boldsymbol{\Phi}\mathbf{h})^{\text{H}}|\mathbf{h}\}+\mathbb{E}\{\mathbf{A}^{\text{H}}\mathbf{n}\mathbf{n}^{\mathrm{H}}\mathbf{A}\})$
@@
Second, the following fact is adopted from [32, Fact 1].
@@
$\boldsymbol{\Phi}\mathbf{x}\overset{\mathrm{d}}{=}\omega\mathbf{x}+\sqrt{\omega-\omega^{2}}\|\mathbf{x}\|\mathbf{W}\alpha$
@@
$-\frac{\sigma_{\text{n}}^{2}(M\bar{N}_{\mathrm{t}}\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i})-N_{\mathrm{t}}(N_{\mathrm{t}}-M)\text{tr}(\boldsymbol{\Omega}_{\mathrm{A}}(\mathbf{x}_{i}))\mathbf{I})}{M(N_{\mathrm{t}}M+N_{\mathrm{t}}-2)}$ (54)
@@
  • [25] P. Sadeghi, P. Rapajic, R. Kennedy, and T. Abhayapala, “Autoregressive time-varying flat-fading channels: Model order and information rate bounds,” in Proc. IEEE Int. Symp. Inf. Theory, 2006, pp. 1061–1065.
@@
  • [33] Remcom, Inc., “Wireless InSite,” 2025, [Online]. Available: https://www.remcom.com/wireless-insite-em-propagation-software/.